Consider a population consisting of the number of teachers per college at small 2-year colleges. Suppose that the number of teachers per college has an average μ-160 and a standard deviation σ 10 (a) Use Chebyshevs Rule to make a statement about the minimum percentage of colleges that have between 130 and 190 teachers. (Round your answer to two decimal places if necessary.) 89.89 (b) Assume that the population is mound-shaped symmetrical. What proportion of colleges have less than 170 teachers? 0.8413
Answer #1

(a)

\mu = 160

\sigma = 10

(130 - 160)/10= - 3

(190 - 160)/10 = 3

By Chebyshev's Inequality:

1-\frac{1}{k^{2}}

of the distribution's values will liewithin k standard deviations of the mean.

Here,

k = 3

Substituting, we get:

1-\frac{1}{k^{2}}=1 - \frac{1}{3^{2}}=0.8889

= 88.89 %

So,

Answer is:

88.89

(b)
To find P(X<170):
Z = (170 - 160)/10 = 1

Table of Area Under Standard Normal Curve gives area = 0.3413

So,

P(X<170) = 0.5 + 0.3413 = 0.8413

So,

Answer is:

0.8413